OpenStudy (anonymous):

Solve 2x2 + x + 7 = 0. Round solutions to the nearest hundredth. x ≈ −4.21 and x ≈ 3.21 x ≈ −2.10 and x ≈ 1.60 x ≈ −1.60 and x ≈ 2.10 No real solutions

OpenStudy (anonymous):

2x^2 + x + 7 = 0 Just use the quadratic equation x = ( -b +- sqrt(b^2-4ac))/2a x = ( -1 += sqrt(-1 - 4*2*7))/4 at this point you can stop because the sqrt is of a negative value meaning there are no real solutions.

OpenStudy (anonymous):

hope i help

OpenStudy (anonymous):

yes u did, can u do another ?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Identify the vertex for the graph of y = −2x2 − 12x + 1. (−3, 19) (−3, 55) (3, −53) (3, −47)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

oo this is easy i did this one

OpenStudy (anonymous):

y=−2x2−12x+1. 0 = -2*X^2-12*X+1 Axis of symmetry: X= -3; Vertex (maximum): ( -3,19) Quadratic formula: X = (-b ±√(b²-4ac))/(2a) = (12±√(144 +8))/( -4) positive discriminant = 152; two real roots: 0.08220700148448801 and -6.082207001484488 factors: (-2X -12.1644)(X -0.0822) hope it help's

OpenStudy (anonymous):

ok which choice would it be ?

OpenStudy (anonymous):

@blackprincesssen